A train of 100 metres travelling at 50 meters per second overtakes another train of length 200 m is moving at 30 m/s. The time taken by the first train to overtake the second is
  1. 5 sec
  2. 10 sec
  3. 15 sec
  4. 20 sec

Grade:11

3 Answers

Nishant
6 years ago
Let the train overtaken covers x distance in t time. Then the train overtaking has to cover x+total length of the trains, i.e., x+300 in same time t.We have two equations in form time = distance/speedOne is t= x/30 which gives x=30t and other is t=300+x/50. Putting value of x we get,t=300+30t/5020t=300t=15 seconds
Sambit
6 years ago
It will take 15 secStepsTotal distance to cover 100+200=300Relative velocity= vel of train A - vel of train B =50-30= 20m/sTime= distance/ velocityTime= 300/20= 15 sec
Nishant
6 years ago
Let the train overtaken covers x distance in t time. Then the train overtaking has to cover x+total length of the trains, i.e., x+300 in same time t.We have two equations in form time = distance/speedOne is t= x/30 which gives x=30t and other is t=300+x/50. Putting value of x we get,t=300+30t/50 20t=300 t=15 seconds